18.01, Calculus, Review Problems, Unit 2

• §2.6: 12, 14

• E9: 1ae, 2 (only parts a,e), 4

• E10: 2, 15, 18

• E11: 4, 6

• §4.6: 5; E13: 1

• E8: 1ae (linear approximations only), 2abc • E31: 1af, 2bf